1. A + 3B → 2C If the rate that B is changing at the beginning of a reactions is 0.546 atm/min, how fast is A changing? Keep the units the same and remember rates are (+) by convention.

Part 2.

A + 3B → 2C

If the rate that B is changing at the beginning of a reactions is 0.521 atm/min, how fast is C changing?

Keep the units the same and remember rates are (+) by convention.
Answer #1

we have relation from eq  

d[A] /dt = (1/3) d[B] /dt , given d[B] /dt = 0.546 atm / min

hence d[A]/dt = ( 1/3) ( 0.546) atm/min = 0.182 atm / min

2) (1/3) d[B] /dt = ( 1/2) d[C] /dt

d[C] /dt = ( 2/3) [d[B] /dt = ( 2/3) ( 0.521 atm/min) = 0.347 atm /min